OpenStudy (anonymous):

An opinion poll asks an SRS of 1500 adults, "Do you happen to jog?" Suppose that the population proportion who jog (a parameter) is p = 0.15. To estimate p, we use the proportion p-hat in the sample who answer "Yes." Justify the use of a normal approximation and find the following probabilities. a) P(p > 0.16) b) P(0.14 < p < 0.16)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

notice how np = 1500*0.15 = 225 and how n(1-p) = 1500(1-0.15) = 1275

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

so because np > 5 and n(1-p) > 5, this allows us to use the normal distribution as an approximation to the binomial distribution given

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

P(p>0.16) :

jimthompson5910 (jim_thompson5910):

so instead of using that binomial distribution calculator from that last problem we did, we use the normal distribution calculator from wolfram alpha

OpenStudy (anonymous):

Kk!

OpenStudy (anonymous):

im there

jimthompson5910 (jim_thompson5910):

ok p = 0.16 means 16%, so take 16% of 1500 to get 0.16*1500 = 240

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

240

jimthompson5910 (jim_thompson5910):

now calculate the mean and std dev mean: 1500*0.15 = 225 std dev: sqrt(1500*0.15*(1-0.15)) = 13.82932

OpenStudy (anonymous):

kk did

jimthompson5910 (jim_thompson5910):

ok now you type in these four values in the four boxes 225 13.82932 240 500 ... (or some other big number 3 std dev away from the mean)

OpenStudy (anonymous):

kk i did

jimthompson5910 (jim_thompson5910):

and you got what

OpenStudy (anonymous):

.139

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

that's your answer to part a

OpenStudy (anonymous):

ok:) now b?

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

14% of 1500 = ??

OpenStudy (anonymous):

210

jimthompson5910 (jim_thompson5910):

so that's your new left endpoint your new right endpoint is 240

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

is the answer 0.7219?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

yay thanks! only a few multiple choice plz then im done i promise ahah